OpenStudy (shadyykatie):

Given ∆ABC with A(0,2), B(-2, -4), and C(4,4). a. Write the equation for the line containing altitude AZ in standard form. b. Write the equation of the line containing midsegment XZ in standard form, where X is the midpoint of AB and Z is the midpoint of BC.

OpenStudy (anonymous):

first find the slope using the given points and using the formula (y2-y1)/(x2-x1) then put it in point slope interecept form y-y1 = m(x-x1) --m is the slope simplify ^^ then put it in standard form standard formula = ax+by = c midpoint = (x1+x2)/2 , (y1,y2)/2

OpenStudy (anonymous):

@ShadyyKatie

OpenStudy (shadyykatie):

im gonna need help I suck at this.

OpenStudy (anonymous):

first you need to B, and find the midpoint for X and Z

OpenStudy (anonymous):

midpoint formula = (x1+x2)/2 , (y1,y2)/2 to find the X use the points A (0, 2), and B (-2, -4) (x1,y1) and (x2,y2) plug them into the midpoint formula what will you have ?

OpenStudy (anonymous):

midpoint formula = (x1+x2)/2 , (y1+y2)/2*

OpenStudy (shadyykatie):

(0 + (-2))/2 (2 + (-4))/2 right?

OpenStudy (anonymous):

yes that's right, now simply that

OpenStudy (anonymous):

(0 + (-2))/2 , (2 + (-4))/2 they are separated by a comma

OpenStudy (anonymous):

(0 + (-2))/2 = -2/2 = -1 (2 + (-4))/2 = -2/2 = -1 so X has Point (-1,-1) now use the mid point formula again to find the Points for Z this time you need to use the points BC as instructed

OpenStudy (shadyykatie):

I got (-1, -1)

OpenStudy (anonymous):

now find Z by using the points BC the same way you found x